Find the force a 67.5kg man would experience while standing on the surface of each of the following planets.

Earth - Mass = 5.98 x 1024 kg - Distance = 6.337 x 105m - Force(N) =?

Mars - Mass = 6.34 x 1023kg - Distance = 3.43 x 105m - Force(N)=?

Pluto - Mass = 5 x 1023kg - Distance = 4 x105m - Force (N) = ?

Show your solutions


1
Expert's answer
2021-12-20T10:27:47-0500


Earth-Mass "F=G\frac{Mm}{r^2}=6.67\cdot10^{-11}\cdot\frac{5.98\cdot10^{24}\cdot67.5}{(6337000)^2}=670.44\ (N)"


Mars-Mass "F=G\frac{Mm}{r^2}=6.67\cdot10^{-11}\cdot\frac{6.34\cdot10^{23}\cdot67.5}{(3430000)^2}=242.6\ (N)"


Pluto-Mass "F=G\frac{Mm}{r^2}=6.67\cdot10^{-11}\cdot\frac{1.3\cdot10^{22}\cdot67.5}{(1188000)^2}=41.47\ (N)"